# Resolution of Simpson’s paradox via the common cause principle

### v1

@@ Appendix A Examples of Simpson’s paradox
Example 7. Here the COVID-19 death rates are compared in Italy and China stanford . According to the data, aggregated death rates in Italy are higher than in China, but in each age group, the death rates are higher in China. Specific details of age groups were not reported in Ref. stanford ; for this example, we take it binary: above 65 years and below it. We assign the death rate – $A_{1}$, country – $A_{2}$, age group – $B$. We propose that the healthcare system can be the common cause $C$ here: a more affordable healthcare system may cause a higher proportion of older people in the country’s society. Indeed, for 2019, Italy had a larger percentage of people aged above 65 than China: 24.05 % versus 12.06 %. On the other hand, the healthcare system will influence death rates in all age groups. When faced with the COVID-19 pandemic, older people have a higher death rate. If the proportion of older people is high, this may result in higher overall death rates, though within each age group, the death rate difference is the opposite.
Appendix B Matrix notations for inverting the common cause equation
Here we develop matrix notations for inverting the common cause equation:
$p(A_{1},A_{2},B)=\sum_{C}p(A_{1},A_{2},C)p(B|C),$ (20)
where the summation goes over all values of $C$. We work for the case when the variables $A_{1}$, $A_{2}$, $B$ and $C$ are binary, though the matrix notations we introduce below are useful more generally.
Eq. (20) can be written in matrix form
$\begin{pmatrix}[ik1]\\ [ik2]\end{pmatrix}=\begin{pmatrix}(1|1)&(1|2)\\ (2|1)&(2|2)\end{pmatrix}\begin{pmatrix}(ik1)\\ (ik2)\end{pmatrix}.$ (27)
where $ik=11,12,21,22$ and the following notations were introduced
$\begin{split}&[111]\equiv p(a_{1},a_{2},b),\hskip 5.0pt[121]\equiv p(a_{1},\bar{a}_{2},b),\dots,\\ &(111)\equiv p(a_{1},a_{2},c),\hskip 5.0pt(121)\equiv p(a_{1},\bar{a}_{2},c),\dots,\\ &(1|1)\equiv p(b|c),\hskip 5.0pt(2|1)\equiv p(\bar{b}|c),\dots,\\ &D=(2|2)+(1|1)-1.\end{split}$ (28)
Inversion of the Eq. (27) gives
$\begin{pmatrix}(ik1)\\ (ik2)\end{pmatrix}=\frac{1}{D}\begin{pmatrix}(2|2)&-(1|2)\\ -(2|1)&(1|1)\end{pmatrix}\begin{pmatrix}[ik1]\\ [ik2]\end{pmatrix}.$ (35)
Eq. (35) implies
$\begin{pmatrix}(i|k1)\{1|k\}\\ (i|k2)\{2|k\}\end{pmatrix}=\frac{1}{D}\begin{pmatrix}(2|2)&-(1|2)\\ -(2|1)&(1|1)\end{pmatrix}\begin{pmatrix}[i|k1][1|k]\\ [i|k2][2|k]\end{pmatrix},$ (42)
where analogously to (28) we introduced the following notations:
$\begin{split}&[1|11]\equiv p(a_{1}|a_{2},b),\hskip 5.0pt[1|21]\equiv p(a_{1},\bar{a}_{2},b),\dots,\\ &(1|11)\equiv p(a_{1}|a_{2},c),\hskip 5.0pt(1|21)\equiv p(a_{1}|\bar{a}_{2},c),\dots,\\ &[1|1]\equiv p(b|a_{2}),\hskip 5.0pt[2|1]\equiv p(\bar{b}|a_{2}),\dots,\\ &\{1|1\}\equiv p(c|a_{2}),\hskip 5.0pt\{2|1\}\equiv p(\bar{c}|a_{2}),\dots.\end{split}$ (43)
The matrix relation (42) results in
$\begin{split}(i|k1)\{1|k\}=\frac{(2|2)}{D}[i|k1][1|k]+\frac{(2|2)-1}{D}[i|k2][2|k].\end{split}$ (44)
Using (44, 35) we get relations employed in the main text:
$(i|k1)$ $=\frac{(2|2)[i|k1][1|k]+((2|2)-1)[i|k2][2|k]}{(2|2)[1|k]+((2|2)-1)[2|k]},$ (45)
$\{1|k\}$ $=\frac{1}{D}(2|2)[1|k]+\frac{1}{D}((2|2))-1)[2|k],$ (46)
$(i|k2)$ $=\frac{((1|1)-1)[i|k1][1|k]+(1|1)[i|k2][2|k]}{((1|1)-1)[1|k]+(1|1)[2|k]},$ (47)
$\{2|k\}$ $=\frac{1}{D}((1|1)-1)[1|k]]+[1|1][2|k]).$ (48)

### v2

@@ 5 Example: COVID-19, Italy versus China
When applying the common cause set-up from section 3.1, we can look at (18), because $A_{2}$ is better described as a label (avoiding dealing with the probability of country). Hence, from the viewpoint of (18), we need a common cause that supplements $A_{2}$ and acts on both $A_{1}$ and $B$. We propose that the quality of healthcare system can be the common cause $C$ here. In particular, a more affordable healthcare system may cause a higher proportion of older people in the country’s society. Indeed, for 2019, Italy had a larger percentage of people aged above 65 than China: 24.05 % versus 12.06 %. On the other hand, the healthcare system will influence death rates in all age groups. If $C$ is binary, then our conclusion from Theorem 1 is opposite to that of [31]: China did worse than Italy.
6 Simpson’s paradox and common cause principle for Gaussian variables
6.1 Formulation of Simpson’s paradox for continuous variables
Simpson’s paradox is uncovered earlier for continuous variables than for the discrete case [1]. Researching the continuous variable paradox and identifying it in big datasets is currently an active research field [23, 24, 29, 47, 48, 49].
The association between continuous variables $A_{1}=\{a_{1}\}$ and $A_{2}=\{a_{2}\}$ can be based on a reasonable definition of correlation coefficient [1, 30]. We focus on Gaussian variables, because this definition is unique for them and amounts to conditional variance. These variables are also important in the context of machine learning (e.g. linear regressions) [50].
Hence the formulation of Simpson’s paradox given $B=\{b\}$ reads instead of (13) [1, 30, 23, 24]:
$\sigma[a_{1},a_{2}]\sigma[a_{1},a_{2}|b]<0~{}~{}{\rm for~{}all~{}}b,$ (41)
$\sigma[a_{1},a_{2}]\equiv\langle(a_{1}-\langle a_{1}\rangle)(a_{2}-\langle a_{2}\rangle)\rangle,$ (42)
$\sigma[a_{1},a_{2}|b]\equiv\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b},$ (43)
$\langle a\rangle_{b}\equiv\int{\rm d}a\,a\,p(a|b),$ (44)
where $\langle a\rangle_{b}$ and $\sigma[a_{1},a_{2}|b]$ are the conditional mean and covariance; $\langle a\rangle$ and $\sigma[a_{1},a_{2}]$ are the mean and covariance; $p(a|b)$ is the conditional probability density of $A=\{a\}$.
The message of (41) is that the usual and conditional covariance have different signs, i.e., they predict different types of associations between $A_{1}$ and $A_{2}$. For instance, $\sigma[a_{1},a_{2}]>0$ means correlation, while $\sigma[a_{1},a_{2}|b]$ implies anti-correlation. Note a subtle difference between this formulation of Simpson’s paradox and that presented in section 2.2. In (4142) the formulation is symmetric with respect to $A_{1}$ and $A_{2}$.
6.2 General solution for Gaussian variables
For fuller generality, we shall assume that $A=\{\boldsymbol{a}\}$, and $B=\{\boldsymbol{b}\}$ are Gaussian column vectors with a number of components (i.e., dimensionality) $n_{A}$, and $n_{B}$, respectively. We also define
$\boldsymbol{y}^{\rm T}=(\boldsymbol{a},\boldsymbol{b}),$ (45)
where ${\rm T}$ means transposition: $\boldsymbol{y}^{\rm T}\boldsymbol{y}$ is a number, while $\boldsymbol{y}\boldsymbol{y}^{\rm T}$ is a matrix. We assume that a Gaussian $n_{X}$-dimensional variable $X=\{\boldsymbol{x}\}$ is the common cause variable for $A$ and $B$:
$P(\boldsymbol{y}|\boldsymbol{x})=\frac{1}{(2\pi)^{n_{A}/2}\,\sqrt{{\rm det}[{\cal Q}]}}\,\,e^{-\frac{1}{2}(\boldsymbol{y}-{\cal C}\boldsymbol{x})^{\rm T}{\cal Q}^{-1}(\boldsymbol{y}-{\cal C}\boldsymbol{x})},$
$P(\boldsymbol{x})=\frac{1}{(2\pi)^{n_{X}/2}\sqrt{{\rm det}[{\cal S}]}}\,\,e^{-\frac{1}{2}\boldsymbol{x}^{\rm T}{\cal S}^{-1}\boldsymbol{x}},$ (47)
${\cal Q}=\begin{pmatrix}{\cal A}&0\\ 0&{\cal B}\end{pmatrix},\qquad\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}={\cal C}\boldsymbol{x},$ (48)
$\langle(\boldsymbol{y}-\langle\boldsymbol{y}\rangle_{\boldsymbol{x}})(\boldsymbol{y}^{\rm T}-\langle\boldsymbol{y}^{\rm T}\rangle_{\boldsymbol{x}})\rangle_{\boldsymbol{x}}={\cal Q},$ (49)
where the common cause feature of $X=\{\boldsymbol{x}\}$ is ensured by the block-diagonal structure of the covariance matrix ${\cal Q}$: ${\cal A}$ and ${\cal B}$ are (resp.) covariance matrices for $A$ and $B$. In (6.2), ${\cal C}$ is $(n_{A}+n_{B})\times n_{X}$ matrix that ensures the coupling between $(A,B)$ and $X$. For simplicity and without loss of generality we assumed that $\langle\boldsymbol{x}\rangle=0$ and hence $\langle\boldsymbol{y}\rangle=0$ in (6.2). We get from (6.2) after arranging similar terms (and omitting normalization):
$P(\boldsymbol{x})P(\boldsymbol{y}|\boldsymbol{x})\propto\exp\bigg{(}-\frac{1}{2}[\,\boldsymbol{x}^{\rm T}-\boldsymbol{y}^{\rm T}{\cal Q}^{-1}{\cal C}V^{-1}\,]$
×$\times\,V\,[\,\boldsymbol{x}-V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}\boldsymbol{y}]$
$-\frac{1}{2}\boldsymbol{y}^{\rm T}[\,{\cal Q}^{-1}-{\cal Q}^{-1}{\cal C}V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}\,]\boldsymbol{y}\,\bigg{)},$ (50)
$V={\cal S}^{-1}+{\cal C}^{\rm T}{\cal Q}^{-1}{\cal C}.$ (51)
Employing (84) we obtain:
${\cal Q}^{-1}-{\cal Q}^{-1}{\cal C}V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}=({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1},$ (52)
$P(\boldsymbol{y})\propto e^{-\frac{1}{2}\boldsymbol{y}^{\rm T}\,({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1}\,\boldsymbol{y}\,},$ (53)
$\langle\boldsymbol{y}\boldsymbol{y}^{\rm T}\rangle={\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T},$ (54)
We now recall (45, 49), introduce the block-diagonal form for ${\cal C}{\cal S}{\cal C}^{\rm T}$, and find
${\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T}=\begin{pmatrix}{\cal A}+{\cal J}&{\cal K}\\ {\cal K}^{\rm T}&{\cal B}+{\cal L}\end{pmatrix},$ (55)
$({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1}=\begin{pmatrix}({\cal A}+{\cal J}-{\cal K}({\cal B}+{\cal L})^{-1}{\cal K}^{\rm T})^{-1}&...\\ ...&...\end{pmatrix},$
where (LABEL:woot) can be deduced via Appendix .4. In that formula we need only the upper-left block, so that all other blocks are omitted. Collecting pertinent expressions from (45, 54, LABEL:woot, 49), we deduce finally
$\langle\boldsymbol{y}\boldsymbol{y}^{\rm T}\rangle={\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T},$ (57)
$\langle(\boldsymbol{y}-\langle\boldsymbol{y}\rangle_{\boldsymbol{x}})(\boldsymbol{y}^{\rm T}-\langle\boldsymbol{y}^{\rm T}\rangle_{\boldsymbol{x}})\rangle_{\boldsymbol{x}}={\cal Q},$ (58)
$\langle(\boldsymbol{a}-\langle\boldsymbol{a}\rangle_{\boldsymbol{b}})(\boldsymbol{a}^{\rm T}-\langle\boldsymbol{a}^{\rm T}\rangle_{\boldsymbol{b}})\rangle_{\boldsymbol{b}}$
$={\cal A}+{\cal J}-{\cal K}({\cal B}+{\cal L})^{-1}{\cal K}^{\rm T}.$ (59)
6.3 The minimal set-up of Simpson’s paradox: 3 scalar variables + scalar cause
For this simplest situation, $\boldsymbol{y}^{\rm T}=(a_{1},a_{2},b)$ is a 3-dimensional vector, ${\cal A}$ is a $2\times 2$ matrix, ${\cal C}$ is a $3\times 1$ matrix, while ${\cal S}$ and ${\cal B}$ are positive scalars. Now (5759) read:
$\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon,$
$0<\epsilon\equiv\frac{{\cal B}}{{\cal B}+{\cal C}^{2}_{31}{\cal S}}<1,$ (60)
$\langle(a_{1}-\langle a_{1}\rangle_{x})(a_{2}-\langle a_{2}\rangle_{x})\rangle_{x}={\cal A}_{12},$ (61)
$\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}.$ (62)
Now consider a scenario of Simpson’s paradox, where
$\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}>0~{}~{}{\rm and}~{}~{}$ (63)
$\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon<0.$ (64)
Due to $10<\epsilon<$ 1, these two inequalities demand ${\cal A}_{12}<0$. Likewise, $\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}<0$ and ${\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon>0$ demand ${\cal A}_{12}>0$. It is seen that under Simpson’s paradox for this minimal situation, the sign of $\langle(a_{1}-\langle a_{1}\rangle_{x})(a_{2}-\langle a_{2}\rangle_{x})\rangle_{x}$ coincides with the sign of $\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}$. We are thus led to the following:
Theorem 2: In the minimal situation (6.362) with the (minimal) common cause, the continuous Simpson’s paradox (41) is resolved in the sense that the decision on the sign of correlations should proceed according to the fine-grained option: $\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}$; see (4142).
For non-minimal common causes, all possibilities of the paradox can be realized; see Appendix .5.
7 Conclusion
We addressed Simpson’s paradox: the problem of setting up an association between two events $a_{1}$, $a_{2}$ given the lurking variable $B$. This decision-making paradox provides two plausible but opposite suggestions for the same situation; see (1) and (2, 3). Either the first option is correct, the second option is correct, or none of them is correct.
We focus on cases when there is a common cause $C$ for $B$ and $A=(A_{1},A_{2})$ (which combines $a_{1}$, $a_{2}$ and their complements). Alternatively, $C$ screens out $A$ from $B$. These cases include those in which there is no causal influence from $A$ to $B$, as well as from $B$ to $A$. Hence, correlations between $A$ and $B$ are to be explained via the common cause $C$, which is a statement of the common cause principle [34, 35]. Now the association between $a_{1}$ and $a_{2}$ is to be decided by looking at $p(a_{1}|a_{2},c)$ for various values of $C$. This task is normally difficult given the fact that $C$ is frequently not fully known and is not observed. However, provided that $A_{1}$, $A_{2}$, $B$ and $C$ are binary, $p(a_{1}|a_{2},c)$ shows the same association as the option (2, 3) of Simpson’s paradox. In this sense, Simpson’s paradox is resolved in the binary situation, provided that the situation allows a binary cause or a binary screening variable. The same conclusion on resolving Simpson’s paradox was reached for Gaussian variables in the minimal situation. Several examples can illustrate the plausibility of a minimal $C$.
These results lead to several interesting research directions. First, in the present paper, we limited ourselves to results that hold for all (minimal) common causes. For many applications this is too stringent: if the common cause is known to exist, but is not observed directly, then it may be sufficient to infer it e.g. via the (generalized) maximum likelihood [42] or the minimal entropy method [41]. This may provide pertinent information on the real common cause and on the structure of Simpson’s paradox. Second, we insisted on a precise common cause. The screening relation (16) is also useful, when it does hold approximately, but the support of $C$ is relatively small. Such an approximate relation (16) provides data compression via feature detection, which is the main message of unsupervised methods such as Non-negative Matrix factorization and Probabilistic Latent Dirichlet indexing [40]. The impact of such approximate, but efficient causes on probabilistic reasoning is an interesting research subject that we plan to explore in the future. Third, the general formalism we developed in section 6 for Gaussian variables may find further applications in the causal analysis of Gaussian machine learning algorithms [50].